OpenStudy (anonymous):

Find the degree measures of the angles whose cosine is -√3 over 2

OpenStudy (anonymous):

\[\cos^{-1} (\frac{ -\sqrt{3} }{ 2 })\]

OpenStudy (linyu):

It's \[\frac{ 2\pi }{ 3 }+2k \pi\]=150+2k180 or\[\frac{ 5\pi }{ 3 }+2k \pi\]=210+2k180 where k=0,1,2,........